A machine costing $212,000 with a four-year life and an estimated $20,000 salvage value is installed in Luther Companys factHaving trouble with the last one
Answer #1

26,500 192,000 20,000 6,500 $ 192,000 Total

Guidance for the last one:

Straight-line rate = (1 / Life years) × 100

                            = 100 / 4

                            = 25%

Double declining (DB) rate = Straight-line rate × 2

                                    = 25% × 2

                                    = 50%

This rate is applicable on the beginning book value in each year (1 to 3) in order to get depreciation expense of that year.

In the 4th year:

Ending book value should be established first; this should be the amount of salvage value, which is $20,000. Once this is established, depreciation expense would be the difference between beginning and ending book values.

Depreciation expense = Ending book value – Beginning book value

                                    = 26,500 – 20,000

                                    = $6,500

The DB rate is not applicable here, since the depreciation expense appears as a balancing figure.

Accumulated depreciation = 185,500 + 6,500

                                                = 192,000.

This is equal to total of depreciation expenses from year 1 to year 4.

Checking:

Accumulated depreciation = Total depreciation expense = Cost – Salvage value = 212,000 – 20,000 = 192,000.